Operations Report – Nov 11th

Crew 215 Operations Report 11-NOV-2019
SOL: 1
Name of person filing report: Shane Usher

Non-nominal systems: Opportunity, Suit 10, Suit 5 Helmet, No power to multiple downstairs wall sockets.

Notes on non-nominal systems: Opportunity brakes do not work, Suit 10 is not functioning (fuse type 5 is blown), Suit 5 Helmet falling apart, Suit 2 Battery supplying low power. Loft Tank Feed tank leaky (temporarily fixed with gaffer tape).
The water pump is very loud.

Generator: run
Hours run: 12h 30 min
From what time last night: 1800
To what time this morning: 0730
List any additional daytime hours when the generator was run: N/A

Solar— SOC % (Before generator is run at night): 78%
Diesel Reading – 54%
Station Propane Reading – 74%
Ethanol Free Gasoline: N/A

Water (loft tank): 55 gallons
Water Meter: 1457225 units
Water (static tank): 304 gallons (not filled)
Static to Loft Pump used – yes
Water in Green Hab: 225 gallons – 28 L (7 gallons) used
Water in Science Dome: 0 gallons

Toilet tank emptied: no

Deimos rover used: not at hab
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –

Spirit rover used: yes
Hours: 110.0
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 100%
Currently charging: yes.

Opportunity rover used: no
Hours: 65.6
Beginning charge: 100%
Ending charge: 100%
Currently charging: no

Curiosity rover used: yes
Hours: 115.7
Beginning charge: (Before EVA) : 100%
Ending charge: (On return from EVA, before charging): 100%
Currently charging: yes

Notes on rovers: no
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used: 0 Gals
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director

HabCar used and why, where? With DIRECTOR
CrewCar used and why, where? At Grand Junction awaiting Crew 216

General notes and comments: N/A

Summary of internet: Functional today, but CAPCOMs was down in the evening on Sol 0.
Summary of suits and radios: Suit 2 needs a replacement battery, Suit 5 Helmet falling
apart, Suit 10 requires type 5 fuse.
Summary of Hab operations: All systems nominal apart from; Stateroom5, Edison screw lightbulb Max 13W required; Fire alarm test button not working in Hab upper level; Fire alarm test button disintegrated in GreenHab; Numerous powerpoints not functional on the ground floor of hab. The water pump is very loud.
Summary of GreenHab operations: Atmospheric equalization to the green hab was carried out during the day, plants were watered (28L). The heater automatically came on at 5 pm.
Summary of Science Dome operations: All systems nominal.
Summary of RAM operations: Not used
Summary of any observatory issues: Nominal
Summary of health and safety issues: All crew healthy

Questions, concerns and requests to Mission Support: Static tank topped up to approximately 350 L. Water heater unusable. Opportunity is currently unusable.
We need advice about how we can obtain a replacement battery and fuse (type 5) for Suits. Loft tank feed pipe needs a new hose connection to eliminate the leak. We might be able to improve the current setup with Teflon tape and duct tape.


Supply Request List:
12-volt lead-acid battery – Suit 2 Hab ground floor
Fuse (Type 5) – Suit 10 Hab ground floor
Fire smoke alarm – Hab upper level
Fire smoke alarm – GreenHab
Water Pump – water tank outside next to Hab?
Water tank feed hose connection – Loft
Gaffer Tape – RAM
Teflon tape – RAM
Electrical Service: Powerpoint not functioning – Hab ground floor near the hot water system
Safety: Static electric shock becoming significant around the Hab

Copyright © The Mars Society. All rights reserved. | Main Site